OpenStudy (anonymous):

Area of a regular decagon with radius of 5 feet? I have the question and answer, but I don't understand how to get the answer. (I'm doing a chapter review test, and we can find the answers, but need to show work.)

geerky42 (geerky42):

Radius of a regular polygon is a distance from center to any vertex: Use this formula to find area:\[A = \dfrac{r^2N \sin \left( \frac{360^o}{N} \right)}{2}\]Where r is radius and N is number of sides.